OpenStudy (anonymous):

-3x(3x-2)+(-x3)(-4x+6)

OpenStudy (whpalmer4):

Are you supposed to simplify that? Could you clarify what "-x3" means? is that supposed to be \(-x^3\) If so, you should use ^ between the x and the 3 to indicate an exponent, or better yet, use the equation editor at the lower left (blue button labeled \(\sum~ \text{Equation}\)

OpenStudy (anonymous):

trying to find the simplest form.. and yes its -x^3

OpenStudy (whpalmer4):

well, what do you get after you apply the distributive property?

OpenStudy (anonymous):

ummm.. I honestly don't know

OpenStudy (whpalmer4):

oh come on. -3x(3x-2) = ?

OpenStudy (whpalmer4):

\[a(b-c) = a*b - a*c\]That's the distributive property

OpenStudy (anonymous):

-9x^2+6?

OpenStudy (whpalmer4):

very good. now how about \[-x^3(-4x+6)=\]?

OpenStudy (anonymous):

4x^4+-6^3

OpenStudy (whpalmer4):

close, you omitted an \(x\) :-)

OpenStudy (anonymous):

but the answer is? 5x^2x^3?

OpenStudy (whpalmer4):

one step at a time. Can you tell me the correct simplification of \(-x^3(-4x+6)\)?

OpenStudy (anonymous):

umm.. 5x^3=6

OpenStudy (anonymous):

+6*

OpenStudy (whpalmer4):

where are you getting that? \[-x^3(-4x+6) = -x^3 * (-4x) -x^3*(6) = \]

OpenStudy (whpalmer4):

\[-x^3(-4x+6) = -x^3 * (-4x) -x^3*(6) \]\[\qquad = 4x^4 - 6x^3\] You had this right before, you just made a mistake in typing it. Then when I asked you to fix the mistake, you started giving me some wacky stuff...

OpenStudy (whpalmer4):

Now you have add the two parts together and collect like terms: \[-3x(3x-2)+(-x^3)(-4x+6) = -9x^2+6x + 4x^4-6x^3\]Usually we write with exponents in descending order, so that would become \[4x^4-6x^3-9x^2+6x\]